# OrionNav: Online Planning for Robot Autonomy with Context-Aware LLM and Open-Vocabulary Semantic Scene Graphs

### v1

@@ Abstract
Enabling robots to autonomously navigate unknown, complex, dynamic environments and perform diverse tasks remains a fundamental challenge in developing robust autonomous physical agents. They must effectively perceive their surroundings while leveraging world knowledge for decision-making. While recent approaches utilize vision-language and large language models for scene understanding and planning, they often rely on offline processing, external computing, or restrictive environmental assumptions. We present a novel framework for efficient and scalable real-time, onboard autonomous navigation that integrates multi-level abstraction in both perception and planning in unknown large-scale environments that change over time. Our system fuses data from multiple onboard sensors for localization and mapping and integrates it with open-vocabulary semantics to generate hierarchical scene graphs. An LLM-based planner leverages these graphs to generate high-level task execution strategies, which guide low-level controllers in safely accomplishing goals. Our framework’s real-time operation enables continuous updates to scene graphs and plans, allowing swift responses to environmental changes and on-the-fly error correction. This is a key advantage over static or rule-based planning systems. We demonstrate our system’s efficacy on a quadruped robot navigating large-scale, dynamic environments, showcasing its adaptability and robustness in diverse scenarios.
@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) has long been the cornerstone for enabling robots to build maps and localize themselves within them [5, 6, 7]. While traditional SLAM techniques primarily focus on geometric and spatial information, they lack the semantic understanding crucial for higher-level reasoning and task execution. Object detection and segmentation pipelines have been previously integrated into these systems to bridge this gap, allowing them to identify and categorize objects within their environment while also refining the map and odometry [8, 9, 10]. Recent advancements in open-vocabulary detection and segmentation [11] are now being integrated into the mapping process, enhancing the robot’s environmental understanding by identifying and classifying objects within varying environments without requiring retraining [12]. However, these approaches often rely on pre-built semantic maps and struggle with real-time operation, limiting their applicability in dynamically changing, large-scale cluttered environments. Effective navigation and task performance in such settings demand scene representations organized hierarchically across multiple abstraction levels. Scene graphs have emerged as a powerful framework to meet this need, capturing objects and their interrelationships in a structured manner [13, 14, 15, 16]. These graphs not only encode geometric information but also capture semantic attributes and relations, providing a concise depiction of the scene to facilitate complex planning using large language models (LLMs) to achieve a desired goal [17, 18, 19, 20, 16].
@@
The advent of LLMs has transformed autonomous planning due to their ability to understand and reason about complex, context-rich scenarios [21, 22, 23, 24, 25, 17, 18, 19, 20, 26]. Thus, our system integrates LLMs with hierarchical scene graph representations, enabling real-time adjustments and error corrections based on the agent’s evolving understanding of its environment. This enables the generation of context-aware high-level plans for autonomous task execution using information from the scene graphs. The execution of these high-level plans is aided by lower-level planners and controllers. We demonstrate this on a quadruped robot, focusing on navigation tasks that involve reaching specific objects as directed by the user, in unknown environments that change over time. This requires robust coordination between high-level planning and lower-level motion control, including obstacle avoidance. This coordination is crucial for effectively translating complex plans into precise physical actions, allowing the robot to navigate and interact with its surroundings in real time.
@@
The quadruped robot utilizes a combination of RGB-D and LiDAR sensors to autonomously navigate an unknown environment by first generating an object map through integrated open-vocabulary semantic segmentation and SLAM. The maps are continuously updated in real time as the robot navigates around the scene. The objects in the map are spatially clustered and associated with specific rooms using an onboard LLM to construct a hierarchical scene graph. This graph representation enables an LLM to formulate a contextually appropriate plan for executing user-defined navigation tasks. A key innovation of our work lies in the system’s ability to continuously update the scene representations and plan online during robot operation. This allows the robot to swiftly respond to environmental changes and correct planning errors on the fly. This capability is crucial for maintaining effectiveness in dynamic, unpredictable real-world settings where static pre-programmed behaviors often fail. The high-level plan guides the low-level motion planner to generate goal-directed navigation commands, which are then refined by an obstacle-avoiding control barrier function (CBF) controller. These refined commands ensure safe and efficient navigation toward the designated goals. The overall flow of the proposed framework is illustrated in Fig. 1. All modules except for the LLM used for planning operate onboard the robot in real-time while demonstrating autonomy in complex environments to perform diverse tasks. Our key contributions are:
@@
  • A framework with multi-level abstractions of perception and planning for autonomous language-driven navigation that adapts to dynamic environments and task failures.
@@
  • A robust and efficient open-vocabulary 3D semantic mapping approach to generate and update hierarchical scene graphs during robot operation.
  • Integration of 3D semantic scene graphs with LLMs for high-level planning that guides lower-level motion planners and CBF-based obstacle avoidance.
@@
  • Implementation of our framework on a quadruped to perform various tasks autonomously in real-world large-scale environments utilizing onboard sensors and computing.
@@ II-A Scene Graph
Scene Graphs have proven to be an efficient and compact way of representing any given environment. This is achieved by using a tree structure with defined nodes representing compressed semantic features derived from high-dimensional map data obtained through Semantic SLAM. Additionally, edges are constructed between node pairs to represent relative spatial or semantic constraints. Such graphs aid in understanding global and local characteristics of the map, which in turn helps with informed and accurate planning. The map data represented by the nodes can be image instances, object instances conveyed through 3D point clouds [27, 13, 28], 2D masks [24], or larger features like annotated spatial regions [29]. To connect these features with respect to each other, edges are coded to represent geometric or semantic relations between the nodes, such as apple <on> table [30]. Such structures have been commonly used for image retrieval [31, 32], object retrieval [16], and task and motion planning [33].
@@
Earlier works have relied on Conditional Random Fields (CRFs) [34] to model graph nodes and edges. Advancements in deep learning have enabled the training of Graph Neural Networks (GNNs) to generate scene graphs from image instances and object masks [35, 36]. However, the advent of Large Visual-Language Models has facilitated the creation of open-vocabulary annotations of nodes and the appending of language-aligned features to improve the performance of downstream tasks. Foundational models [37, 38, 39, 40] are widely used to enrich scene graph nodes with semantic features such as CLIP [37] embeddings and linguistic annotations like labels, captions, and descriptions. Many methods [13, 14, 15, 16] associate object nodes with corresponding CLIP features obtained by fusing individual embeddings from multiple 2D views, masked views, or re-projected 3D features. Works like ConceptGraphs augment the nodes with language captions to enable the use of generated scene graphs with LLM-based planners. Along with node enhancement, foundational models are also capable of generating semantic relations between nodes. VLMs such as LLaVa [39] or BLIP [41] are used either to directly generate spatial semantic relations between objects using SLAM-derived poses and 2D camera views [24], or indirectly to train graph networks to populate the node edges [13]. However, most of these approaches are computationally intensive, preventing online scene graph generation. This limits the applicability of such works to static environments. Our method, on the other hand, realizes a novel scene graph representation and generation, enabling it to work online.
@@
The works closest to our approach are HOV-SG [16] and Clio [15], which is an extension of their previous closed-set scene graph generation method called Hydra [42]. HOV-SG segments the scene using the Segment Anything Model [43], projects the segments as point clouds, fuses CLIP-segment features, and clusters them to generate 3D masks. These are used to generate a scene graph for indoor environments [44, 45], where nodes are characterized using fused CLIP features and heuristics to generate a scene graph with a hierarchy of buildings, floors, and rooms. However, their computationally expensive clustering and classification methods make it difficult to implement in real time. On the other hand, Clio attempts efficient graph generation using the Information Bottleneck framework. By clustering the scene into task-conditioned primitives, they limit the scene graph to task-relevant semantic concepts, leading to efficient scene graph generation in real time for object retrieval. However, such scene graphs are applicable only to a predefined task list, preventing usage on out-of-list tasks. Our method aims to mitigate the shortcomings of both these methods while enabling real-time scene graph generation and execution of various tasks.
@@ II-B Semantic Mapping
Scene segmentation is a fundamental task in computer vision that aims to partition an image or a 3D scene into semantically meaningful regions, playing a crucial role in applications such as object recognition, scene understanding, and robot navigation [46, 47, 2, 4, 48, 49, 50]. Deep learning has revolutionized this field, enabling the development of highly accurate and efficient models across various paradigms. These include semantic segmentation (inferring high-level category concepts), instance segmentation (grouping foreground pixels into distinct object instances), panoptic segmentation (combining semantic and instance segmentation), and entity segmentation (inferring masks of object instances without classification) [51, 52, 11]. Panoptic segmentation, in particular, predicts both class and instance labels for each pixel, assigning instance labels to foreground objects (‘things’) and single labels to background parts (‘stuff’). The advent of foundation vision-language models like CLIP [37] has enabled more sophisticated approaches to fine-grained recognition and segmentation of objects in scenes in an open vocabulary setting [12]. These models leverage large-scale pre-training on diverse image-text pairs, allowing them to generalize to a wide range of visual concepts and enabling zero-shot learning capabilities which are made robust with out-of-distribution detection [53]. Recent developments have further extended these capabilities to open-vocabulary semantic and instance segmentation tasks, demonstrating impressive zero-shot generalization to novel object categories [11, 54, 55, 56, 57, 58, 59, 60].
@@
Reconstruction of 3D structure with object recognition and grouping remains a fundamental challenge in computer vision, especially in the context of 3D scene understanding. These methods help a large language model help understand the scene through reasoning [61, 62] for building grounded 3D large language models [63, 64, 65, 66, 67, 68, 69]. The semantic information aids in both understanding the scene as well as improving the accuracy of the map and odometry for semantic SLAM [8, 70, 71, 72, 73, 74, 75, 76, 77, 78]. End-to-end 3D methods that predict 3D semantic information directly from sensor modalities like camera and depth, or from intermediate representations like reconstructed point clouds or voxel-based representations, can capture complex 3D relationships. Recent works have significantly improved the performance of open vocabulary 3D segmentation using point clouds from LiDAR or depth data from RGB-D cameras [79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. In addition, there have also been approaches fusing embeddings from open world vision language models like CLIP with neural radian fields, Gaussian splats or implicit neural network based representation [29, 89, 90, 91] for generating a consistent segmentation of the scene. The open vocabulary allows for greater flexibility in real-world applications and can help an LLM to understand the scene with multiple objects, which multi-modal LLMs hallucinate or fail to perceive [92]. However, these methods do not scale well in larger scene making them computationally intensive and memory-consuming, rendering them impractical for real-time segmentation in robotic applications.
@@
OrionNav draws inspiration from prior works while prioritizing efficiency, striking a balance between adaptability and real-time performance. The proposed method leverages the strengths of both 2D and 3D approaches, utilizing the efficiency and generalization capabilities of 2D CLIP-based models while incorporating 3D geometric information to enhance spatial understanding. It integrates a CLIP-based open vocabulary 2D segmentation model with 3D geometric information from map and odometry for improved scene understanding. We utilize the FC-CLIP [93] model with a ConvNeXt [94] CLIP [37] backbone, enabling efficient processing of high-resolution images where transformer-based models struggle. This architecture maintains translation equivariance and allows multi-scale feature fusion, crucial for detailed environmental segmentation in robotic applications. The system performs lightweight updates to the semantic object map leveraging spatial and semantic properties of the scene and observed objects in it. This map can be further utilized by a scene graph to generate hierarchical object clusters and relationships. This hybrid approach offers a robust, efficient solution that addresses trade-offs between computational efficiency and comprehensive 3D scene understanding.
@@ II-C Multi-modal Language Models for Robotic Task Planning
Robotic task planning research, in recent years, has benefitted from the use of LLMs [21, 22, 23, 25, 17, 18, 19, 20, 26]. Early studies in this field primarily focused on manipulation tasks and humanoid simulations [21, 22, 23]. Recent works have shifted toward generating plans for mobile robots operating in larger environments [17, 18, 19, 20, 26, 95]. Strategies such as environment feedback through the use of the Python assert keyword in the plans [21], affordance functions [25], and success detection and scene description [23] have been used to ground the LLM planners in the environment. Studies focusing on mobile robots operating in large environments often utilize scene graphs to ground LLMs in their environment [17, 18, 19, 20, 24, 16], a strategy we also adopt in this framework.
ConceptGraphs [24] and HOV-SG [16] use LLM to retrieve the target object from the prebuilt scene graph while we utilize an LLM planner that acts as a high-level controller and participates in the hierarchical scene graph generation process with its commands. While SayPlan [17], MoMa-LLM [20], and SayNav [18] use LLMs and scene graphs in a similar way to our approach, our method distinguishes itself through broader assumptions and a more comprehensive experimental design. SayPlan [17] assumes access to a prebuilt scene graph, whereas we iteratively update the scene graph as the robot explores the environment. MoMa-LLM [20] assumes access to ground-truth perception data, including semantic masks, depth, localization, and handle detection. Similarly, SayNav [18] relies on known robot pose and semantic masks for objects. Although these methods demonstrate real-world robot applications, their reliance on such assumptions limits their deployment in environments where those conditions are unmet. To address this limitation, we develop an end-to-end system that does not rely on these assumptions and integrates our LLM Planner within it.
Some works incorporate different techniques instead of exploiting LLMs with scene graphs for robot navigation tasks. For example, Explore until Confident [96] employ Vision-Language Models (VLMs) to build a semantic map for embodied question answering, replacing frontier-based search by leveraging semantic knowledge of the environment. OK-Robot [97] proposes an end-to-end system for open-ended pick-and-drop tasks using pretrained models. They employ CLIP [37] to convert language queries into semantic vectors, which are then used to search for the closest voxel in a semantic point cloud. PixelNav [95] proposes a direct RGB-based solution that utilizes an LLM planner for waypoint generation in RGB space for object navigation tasks. However, most of these methods either do not create an environment representation that can be used for subsequent tasks in the same environment, or require a prebuilt scene representation.

### v2

@@ Abstract
Enabling robots to autonomously navigate unknown, complex, dynamic environments and perform diverse tasks remains a fundamental challenge in developing robust autonomous physical agents. These agents must effectively perceive their surroundings while leveraging world knowledge for decision-making. Although recent approaches utilize vision-language and large language models for scene understanding and planning, they often rely on offline processing, offboard compute, make simplifying assumptions about the environment and perception, limiting real-world applicability. We present a novel framework for real-time onboard autonomous navigation in unknown environments that change over time by integrating multi-level abstraction in both perception and planning pipelines. Our system fuses data from multiple onboard sensors for localization and mapping and integrates it with open-vocabulary semantics to generate hierarchical scene graphs from continuously updated semantic object map. The LLM-based planner uses these graphs to create multi-step plans that guide low-level controllers in executing navigation tasks specified in natural language. The system’s real-time operation enables the LLM to adjust its plans based on updates to the scene graph and task execution status, ensuring continuous adaptation to new situations or when the current plan cannot accomplish the task, a key advantage over static or rule-based systems. We demonstrate our system’s efficacy on a quadruped navigating dynamic environments, showcasing its adaptability and robustness in diverse scenarios. Project video: https://www.youtube.com/watch?v=gqWCSjLLG1E.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) has long been the cornerstone for enabling robots to build maps and localize themselves within them [5, 6, 7]. While traditional SLAM techniques primarily focus on geometric and spatial information, they lack the semantic understanding that is essential for higher-level reasoning and task execution. Object detection and segmentation pipelines have been previously integrated into these systems to bridge this gap, allowing them to identify and categorize objects within their environment while also refining the map and odometry [8, 9, 10]. Recent advancements in open-vocabulary detection and segmentation [11] are now being integrated into the mapping process, enhancing the robot’s environmental understanding by identifying and classifying objects within varying environments without requiring retraining [12]. However, these approaches often rely on pre-built semantic maps and struggle with real-time operation, limiting their applicability in dynamically changing, large-scale cluttered environments. Effective navigation and task performance in such settings demand scene representations organized hierarchically across multiple abstraction levels. Scene graphs have emerged as a powerful framework to meet this need, capturing objects and their interrelationships in a structured manner [13, 14, 15, 16]. These graphs not only encode geometric information but also capture semantic attributes and relations, providing a concise depiction of the scene to facilitate complex planning using large language models (LLMs) to achieve a desired goal [17, 18, 19, 20, 16].
@@
The advent of LLMs has transformed autonomous planning due to their ability to understand and reason about complex, context-rich scenarios [21, 22, 23, 24, 25, 17, 18, 19, 20, 26]. Thus, our system integrates LLMs with hierarchical scene graph representations, enabling real-time adjustments and error corrections based on the agent’s evolving understanding of its environment. This enables the generation of context-aware high-level plans for autonomous task execution using information from the scene graphs. The execution of these high-level plans is aided by lower-level planners and controllers. We demonstrate this on a quadruped robot, focusing on navigation tasks that involve reaching specific objects as directed by the user, in unknown environments that change over time. This requires robust coordination between high-level planning and lower-level motion control, including obstacle avoidance. This coordination is crucial for effectively translating simple natural language instructions into complex plans and then into precise physical actions, enabling the robot to navigate autonomously to specific objects or rooms based on conversational commands.
@@
Our proposed OrionNav framework utilizes a quadruped robot equipped with a combination of RGB-D and LiDAR sensors to autonomously navigate an unknown environment to execute a remote user’s command. It generates an object map through integrated open-vocabulary semantic segmentation and SLAM that are continuously updated in real time as the robot navigates around the scene. The objects in the map are spatially clustered and associated with specific rooms using an onboard LLM to construct a hierarchical scene graph. This graph representation enables another LLM to formulate a contextually appropriate plan for executing user-defined navigation tasks. A key innovation of OrionNav lies in the system’s ability to continuously update the scene representations and plan online during robot operation. This allows the robot to swiftly respond to environmental changes and correct planning errors on the fly. This capability is essential for maintaining effectiveness in dynamic, unpredictable real-world settings where static pre-programmed behaviors often fail. The high-level plan guides the low-level motion planner to generate goal-directed navigation commands, while an obstacle-avoiding control barrier function (CBF) controller reactively adjusts the robot’s path when obstacles are detected. These adjustments ensure safe and efficient navigation toward the designated goals. The overall flow of the proposed OrionNav framework is illustrated in Fig. 1. All modules except for the LLM used for planning operate onboard the robot in real-time while demonstrating autonomy in complex environments to perform diverse tasks. Our key contributions are:
@@
  • A real-time, language-driven autonomous navigation that employs multi-level abstractions of perception and planning to execute diverse tasks while handling changes in environments and recovery from task failure.
@@
  • A robust and efficient open-vocabulary 3D semantic mapping approach for real-time generation and updating of hierarchical scene graphs during robot operation.
  • A LLM-based planner that uses hierarchical scene graphs to create multi-step plans for navigation tasks with ability to adjust plan based on environmental changes, task progress, and potential task failure.
  • An enhanced exploration strategy integrated with low-level motion planners and CBF-based controller to enable a robot to operate safely in dynamically changing environments while executing diverse navigation tasks.
@@
  • Real-world implementation of the framework on a quadruped robot, autonomously performing various real-time navigation and search tasks in environments using onboard sensors and computing resources.
@@ II-A Scene Graph
Scene Graphs have emerged as an efficient and compact method for representing environments. They employ a tree structure consisting of ‘nodes’ that represent environmental features, connected by ‘edges’ that denote relative spatial or semantic relationships between node pairs. These graphs facilitate in understanding both global and local characteristics of the environment, thereby enabling more informed and accurate planning for executing a given task. The scene information represented by the nodes can be image and object instances conveyed through 3D point clouds [27, 13, 28], 2D masks [24, 18, 20], or language-guided 3D representations using radiance fields [29, 30, 31, 32]. Edges are encoded to represent geometric or semantic connections between nodes, establishing relationships between these features. For example, an edge might indicate that an apple is on a table [33]. Such structures have been commonly used for image retrieval [34, 35], object retrieval [16], and task and motion planning [36].
@@
Traditionally, Conditional Random Fields [37] were used to model graph nodes and edges but advancements in deep learning have enabled the use of graph neural networks to generate scene graphs from image instances and object masks [38, 39]. The advent of Large Visual-Language Models has facilitated the creation of open-vocabulary annotations of nodes and the use of language-aligned features to improve performance on downstream tasks. Foundational models [40, 41, 42, 43] have been used to enrich scene graph nodes with semantic features like Contrastive Language-Image Pre-training (CLIP) [40] embeddings and linguistic annotations like labels, captions, and descriptions. Next, object nodes associated with corresponding CLIP embeddings obtained by fusing multiple 2D views, masked views, or re-projected 3D features [13, 14, 15, 16]. However, these approaches are typically computationally expensive, preventing online scene graph generation. On the other hand, OrionNav uses a lightweight scene graph generation process, enabling it to work online. ConceptGraphs [24] augment the nodes with language captions to enable the use of generated scene graphs with LLM-based planners. Along with node enhancement, foundational models are also capable of generating semantic relations between nodes. Large Vision-Language models (VLMs) and LLMs such as LLaVa [42], BLIP [44] or GPT [41] are used either to directly generate spatial semantic relations between objects using SLAM-derived poses and 2D camera views [24], or indirectly to train graph networks to populate the nodes and edges [13]. Within the domain of scene graph generation, two approaches closely align with our work: HOV-SG [16] and Clio [15].
@@
HOV-SG [16] obtains 2D objects masks from image instances using the Segment Anything Model [45]. Their fused CLIP features are aggregated for 3D point corresponding to the objects, and later processed to identify a single feature for each object node. These are added to a hierarchical scene graph [46, 47], where heuristics are used to segment floors and rooms and their corresponding nodes are characterized using fused CLIP features. However, their computationally expensive clustering and classification methods make it difficult to implement them in real time. On the other hand, Clio [15] extends the Hydra framework [48] to generate scene graphs with open vocabulary labels, while introducing a novel approach to determine appropriate object representation granularity. It employs an Information Bottleneck formulation to cluster 3D primitives into task-relevant semantic concepts, thus enhancing the scene graph’s utility for specific tasks while generating it efficiently. However, such scene graphs are constrained by their reliance on predefined task sets, limiting their adaptability to unforeseen scenarios and tasks. Our method mitigates these shortcomings by facilitating real-time scene graph generation and allowing the execution of a diverse range of tasks, including those not originally anticipated.
@@ II-B Semantic Mapping
Scene graph generation relies on information about the environment acquired through mapping and semantic labeling process. These processes integrate image or point cloud segmentation techniques with spatial mapping data generated from SLAM frameworks. SLAM provides crucial geometric and structural information about the environment, complementing the semantic understanding derived from segmentation. This synergy enables the creation of robust and contextually rich representations through semantic object maps, enabling better scene understanding for autonomous robot navigation and decision-making. [49, 50, 2, 4, 51, 52, 53]. Segmentation techniques are fundamental to semantic labeling and scene graph generation, enabling detailed parsing of visual information to improve the quality and detail of scene representations. The field encompasses various approaches, each contributing unique capabilities to the interpretation of complex visual scenes. Semantic segmentation infers high-level category concepts, while instance segmentation groups foreground pixels into distinct object instances. Panoptic segmentation combines these two approaches, and entity segmentation infers masks of object instances without classification [54, 55, 11]. Panoptic segmentation, in particular, predicts both class and instance labels for each pixel, assigning instance labels to foreground objects (‘things’) and single labels to background parts (‘stuff’). The advent of foundation VLMs like CLIP [40] has enabled more sophisticated approaches to fine-grained recognition and segmentation of objects in scenes in an open vocabulary setting [12]. These models leverage large-scale pre-training on diverse image-text pairs, allowing them to generalize to a wide range of visual concepts and enabling zero-shot learning capabilities which are made robust with out-of-distribution detection [56]. Recent developments have further extended these capabilities to open-vocabulary semantic and instance segmentation tasks, demonstrating impressive zero-shot generalization to novel object categories [11, 57, 58, 59, 60, 61, 62, 63].
@@
Semantic mapping techniques, vital for scene graph generation, extend beyond 2D segmentation to encompass complex 3D structures. Reconstruction of 3D structures with object recognition and grouping remains a fundamental challenge in computer vision, especially in the context of 3D scene understanding. These methods help a large language model understand the scene through reasoning [64, 65] for building grounded 3D large language models [66, 67, 68, 69, 70, 71, 72]. The semantic information aids in both understanding the scene as well as improving the accuracy of the map and odometry for semantic SLAM [8, 73, 74, 75, 76, 77, 78, 79, 80, 81]. End-to-end 3D methods that predict 3D semantic information directly from sensor modalities like camera and depth, or from intermediate representations like reconstructed point clouds or voxel-based representations, can capture complex 3D relationships. Recent works have significantly improved the performance of open vocabulary 3D segmentation using point clouds from LiDAR or depth data from RGB-D cameras [82, 83, 84, 85, 86, 87, 88, 89, 90]. In addition, there have also been approaches fusing embeddings from open world VLMs like CLIP with neural radian fields, Gaussian splats or implicit neural network based representation [29, 30, 31, 32] for generating a consistent segmentation of the scene. The open vocabulary allows for greater flexibility in real-world applications and can help an LLM to understand the scene with multiple objects, which multi-modal LLMs hallucinate or fail to perceive [91]. However, these methods do not scale well in larger scenes making them computationally intensive and memory-consuming, rendering them impractical for real-time segmentation in robotic applications.
@@
OrionNav draws inspiration from prior works while prioritizing efficiency, striking a balance between adaptability and real-time performance. The proposed method leverages the strengths of both 2D and 3D approaches, utilizing the efficiency and generalization capabilities of 2D CLIP-based models while incorporating 3D geometric information to enhance spatial understanding. It integrates a CLIP-based open vocabulary 2D segmentation model with 3D geometric information from maps and odometry for improved scene understanding. We utilize the FC-CLIP [92] model with a ConvNeXt [93] CLIP [40] backbone, enabling efficient processing of high-resolution images where transformer-based models struggle. This architecture maintains translation equivariance and allows multi-scale feature fusion, crucial for detailed environmental segmentation in robotic applications. The system performs lightweight updates to the semantic object map leveraging spatial and semantic properties of the scene and observed objects in it. This map can be further utilized by a scene graph to generate hierarchical object clusters and relationships. This hybrid approach offers a robust, efficient solution that addresses trade-offs between computational efficiency and comprehensive 3D scene understanding.
@@ II-C Multi-Modal Language Models for Robotic Task Planning
In recent years, robotic task planning has benefitted from the use of LLMs [21, 22, 23, 25, 17, 18, 19, 20, 26]. Early studies in this field primarily focused on manipulation tasks and humanoid simulations [21, 22, 23]. Recent works have shifted toward generating plans for mobile robots operating in larger environments [17, 18, 19, 20, 26, 94]. Strategies such as environment feedback through the use of the Python assert keyword in the plans [21], affordance functions [25], and success detection and scene description [23] have been used to ground the LLM planners in the environment. In addition, robot navigation frameworks have also employed VLMs to build semantic maps for embodied question answering, replacing frontier-based search by leveraging semantic knowledge of the environment. OK-Robot [95] proposed an end-to-end system for open-ended pick-and-place tasks using pre-trained models. Specifically, they utilize CLIP [40] to convert language queries into semantic vectors, which is then used to search for the semantically closest voxel in a semantic point cloud. PixelNav [94] proposes a direct RGB-based solution that utilizes an LLM planner for waypoint generation in RGB space for object navigation tasks. However, these methods either create an environment representation that cannot be used for subsequent tasks in the same environment or require a prebuilt scene representation.
Studies focusing on mobile robots operating in large environments have often utilized scene graphs to ground LLMs in their environment [17, 18, 19, 20, 24, 16], a strategy we also adopt in OrionNav. ConceptGraphs [24], while focusing primarily on scene graph generation, employed an LLM-based planner to retrieve target objects from a pre-built scene graph as a downstream task. Similarly, HOV-SG [16] emphasized scene graph generation and utilized an LLM to parse complex natural language queries for object retrieval from the pre-built scene graph. In contrast, we introduce an LLM planner with several action primitives that function as a high-level controller, actively participating in the scene graph generation process. While SayPlan [17], MoMa-LLM [20], and SayNav [18] also leverage LLMs as high-level planners, they face constraints that limit their practical application. SayPlan [17] requires pre-built scene graphs, whereas MoMa-LLM [20] and SayNav [18] rely on known robot poses and semantic object masks. In contrast, OrionNav overcomes these constraints by performing perception, planning, and control for navigation tasks using onboard sensors in real-time, enabling autonomous navigation in diverse, unknown, and dynamic real-world environments.